Blue Mountain Corp. has a share price of $30 today. If the company is expected to pay a dividend of $0.60 this year, and its stock price is expected to grow to $34.5 at the end of the year. Calculate the dividend yield and the capital gain yield respectively. HTML Editor Keyboard Shortcuts
Answer #1
Dividend yield = 0.60/30 = 2.00%
Capital gains yield = 34.5/30-1 = 15.00%
Total yield 17.00%